OpenStudy (anonymous):

Solve the problem. A can in the shape of a right circular cylinder is required to have a volume of 700 cubic centimeters. The top and bottom are made up of a material that costs 12¢ per square centimeter, while the sides are made of material that costs 5¢ per square centimeter. Find a function that describes the total cost of the material as a function of the radius r of the cylinder.

OpenStudy (anonymous):

well first we know that this is going to be an area problem... the cost is going to depend on the area of material used... so do you know the area of a right circular cyl? and also you have to add the top and bottom, so those would be circles...so what are the equations for area of a cyl. and a circle?

OpenStudy (anonymous):

C(r) = 0.24πr^2 + (70/r) C(r) = 0.08πr^2 + (140/r) C(r) = 0.16πr^2 + (140/r) C(r) = 0.08πr^2 + (70/r)